# CodeMind: A Framework to Challenge Large Language Models for Code Reasoning

### 15 Feb 2024 v1

@@ Abstract
Our extensive evaluation of nine LLMs across five benchmarks in two different programming languages using CodeMind shows that LLMs fairly understand control flow constructs and, in general, are capable of reasoning how inputs evolve to output, specifically for simple programs and the ones they can correctly synthesize. However, their performance drops for code with higher complexity, non-trivial logical and arithmetic operators, non-primitive types, and API calls. Furthermore, we observe that, while correlated, specification reasoning (essential for code synthesis) does not imply execution reasoning (essential for broader programming tasks such as testing and debugging): ranking LLMs based on test passing can be different compared to code reasoning.
@@ 1 Introduction
Using CodeMind, we performed a large-scale ground theory study to assess LLMs for code reasoning. We selected nine models, including both general-purpose and Code LLMs and prompted them for IER, DER, and SR tasks on 5395 programs written in Java and Python [^1] These programs are from five programming benchmarks, namely HumanEval (Chen et al., 2021), MBPP (Odena et al., 2021), CRUXEval (Gu et al., 2024) CodeNet (Puri et al., 2021), and Avatar (Ahmad et al., 2021).. We observe that:
@@
Our contributions are: (1) CodeMind framework for code reasoning that formally defines three inductive code reasoning tasks. CodeMind is open-source (CodeMind, 2024) and accepts contributions from researchers to integrate more code reasoning tasks into it. (2) A large-scale ground-theory evaluation of LLMs for code reasoning using CodeMind. (3) A comprehensive, in-depth analysis of results that offers a catalog of root causes negatively impacting the abilities of LLMs for code reasoning. This catalog would be a valuable guideline for developing better benchmarks that truly evaluate the programming abilities of LLMs.
@@ 2 Related Work
Zhang et al. (2024) investigate transformers’ ability to learn or infer the recursive patterns from input and output pairs. They conclude that due to the inherent limitations of transformers, they may fail to learn recursion and instead find shortcut algorithms to reason about how outputs are related to inputs. Compared to this work, we evaluate LLMs regardless of architecture and training data but from the program perspective. We show LLMs can understand recursion but usually lose track of data flow due to the inability to correctly reason about loop conditions.
@@ 3 CodeMind
Program specification defines a function $S:S_{I}\rightarrow S_{O}$, where $S_{I}$ is a set of all possible inputs to the program and $S_{O}$ is a set of corresponding outputs. A code synthesized based on the implementation is usually a (partial [^2] Due to the specification’s abstraction and ambiguities, especially when given in natural language, implementation is a partial function of what specification defines.) function $C:C_{I}\rightarrow C_{O}$. We define a program to be correct with respect to specification, if it satisfies all the following conditions:
@@ 3.1 Execution Reasoning
IER evaluates LLMs for any arbitrary code for general inductive code reasoning, which requires understanding code constructs, arithmetic and logic operations, and control flow. However, even for human developers, reasoning about their developed code is easier than any arbitrary code. Furthermore, as a self-consistency (Min et al., 2023) measurement, LLMs should be able to reason about the code they can correctly synthesize. This demands to have the following execution reasoning task.
@@ 5 LLM Evaluation on IER
  • Moving down in the table, the models face greater challenges in IER, i.e., reasoning about the execution on CodeNet and Avatar programs, compared to MBPP, HumanEval, and CRUXEval. One potential reason is the complexity of such programs as demonstrated in Figure 2. A detailed breakdown of the model’s performance (Figure 3) shows a strong negative correlation [^3] Spearman’s Rank Order Correlation (ROC) (Spearman, 1961) between CC and CRR. between CRR and CC, confirming that models struggle more in IER for a more complex code. At the same time, some models, namely Llama 2, CodeLlama, MagicCoder, StarCoder, and WizardCoder, achieve a lower performance on CRUXEval compared to HumanEval, which are less complex regarding both LoC and CC. This entails a further better understanding of what factors other than CC impact the CRR performance of the models (section 8).
@@ 6 LLM Evaluation on DER
We seek to address the critical question of how effectively the model can correctly reason about the correct programs it has generated. This evaluation requires us to align code synthesis and code reasoning tasks together. Our pipeline for evaluating DER consists of three steps: (1) following the best practices, we prompted subject LLMs for code synthesis; (2) we ran the synthesized program against existing tests; and (3) for the programs with test pass, we prompted the model for code execution reasoning using the chosen test input and under CoT style. Note that we also removed the comments from the synthesized code for fairness. We excluded the programs from CRUXEval, CodeNet, and Avatar, since these datasets are not designed for code synthesis and lack proper program specifications. Also, we could not reproduce the code synthesis results of Llama 2 and excluded that from subject LLMs. Similar to IER experiments, we set the temperature to zero  [^4] As a result of this design decision, our synthesis results might be different from existing leaderboards. to account for the non-determinism and reproducibility of the results.
@@ 8 In-Depth Analysis of Results
We further analyzed the IER results, which evaluate the general ability of LLMs in code reasoning. In the first step, we wanted to see if LLMs understand different code constructs. Without knowing the logic of each code construct, reasoning about code execution is impossible. To that end, we tagged each of $5395$ programs based on code constructs used in their implementation with the following labels: For, While, If, Try, Switch, Nested Loop, Nested If, and Basic. A program tagged with a Basic label has no special code construct. Next, we clustered the programs per tag and computed the CRR of LLMs for each cluster. Figure 5 shows the results of this analysis for the top five best-performing LLMs. We can observe that models handle conditional statements better than recursion, except for Try-Catch or Try-Except statements. Furthermore, when it comes to nested constructs, the CRR values notably drop. Impact of Loop Properties. Given that models struggle the most with recurring constructs, we focused the programs with For, While, and Nested Loop tags at the next step. We hypothesize this struggle is due to the loop’s length or determining the loop’s length. The former questions whether it is harder for the model to track the program’s data flow as loops get longer. The latter questions the abilities of models to reason about how many times a code block should be repeated, regardless of how long the length would be.
Figure 6 plots the distribution of correct versus incorrect cases and CRR values per each loop length in Java programs. Sub-figure labels show the ROC coefficients between the loop length and CRR. We can observe a moderate to strong negative correlation between the loop length and CRR of the models, i.e., CRR decreases as the loop length increases. By manually investigating the incorrect IER cases, we also noticed that LLMs mostly failed to reason about loop conditions correctly. Without understanding loop conditions and understanding the number of iterations, it is impossible to reason about the execution correctly. Furthermore, we found cases where, although the model could reason about loop conditions and the number of iterations, it lost track of data flow in the loop, thereby predicting output incorrectly.
@@
All models failed in IER of this relatively simple program. Looking at the CoT analysis, GPT-3.5 correctly understands the loop condition and can reason about all the variable values in the first and second iterations. However, it stays in the loop for one more. StarCoder cannot correctly reason about loop condition and iterates the code block six times. MagicCoder CoT only explains the code (correct explanation) without communicating how it reasons about variables.
@@ 9. In third iteration, amount becomes 15 and time
Impact of if-else Properties. Our manual investigations reveal that LLMs struggle with conditional constructs for two reasons: (1) confusion about branches and (2) wrong reasoning about conditional predicates. Concerning the former, we observed that LLMs sometimes struggle to understand the meaning of conditional statements. For example, regardless of which branch to take, they always go into both “if” and “else” branches to reason about execution. However, such cases rarely happen, and in most cases, LLMs struggle to analyze the conditional predicates correctly, take the wrong branch, and predict the output incorrectly.
Impact of Math and Logic Operators. Understanding the meaning of math and logic operators is essential to reasoning how inputs evolve into output through execution. The observations from previous analyses also underscore their importance to reason about the program control flow. Our manual investigation of failed IER cases shows the prevalence of arithmetic and logic operators in programs, with the former dominating the number of the latter. LLMs struggled more to reason about bitwise operators (e.g., $>$, $\wedge$, >>, <<, |, and &\&&) compared to arithmetic (e.g., + and -) and comparison operators (e.g., >, <).
@@
StarCoder, while failing to predict the output, correctly understands the meaning of most operators. For example, it can reason that the bitwise AND of n and n+1 is used to determine whether integer n is a power of 2. It can also reason that the loop condition checks if the least significant bit of n is set to 0. However, it seemingly cannot understand the meaning of the return statement (setting the least significant bit to 1), hence failing to predict the correct output.
@@
Tuples and Lists consist of multiple items and every single one of them may change during the program execution. As a result, it is unsurprising that models struggle to track the flow of inputs through potentially different execution paths and reason about a complex output as a whole. Additionally, given that manipulation of such types involves API calls, e.g., min(), next(), charAt(), understanding changes requires additional efforts by LLMs.
@@ 9 Concluding Remarks
In this paper, we discussed the necessity of code reasoning tasks as an alternative to evaluate LLMs for programming tasks. We introduced CodeMind, a framework that supports several code reasoning tasks, and used CodeMind in a large-scale grounded theory study to evaluate state-of-the-art LLMs for code reasoning. Our results demonstrate that LLMs, in general, understand code constructs and are capable of reasoning about program specification and following how inputs evolve to output through execution. However, their ability is limited as the code becomes more complex, i.e., has more complex control- or data flow, contains non-primitive types, and invokes API calls. We also observe that specification reasoning, which is essential to generate a code from a given program specification, does not mean models can also reason about code execution. We are considering two future directions based on this work. First, we plan to add more code reasoning tasks to CodeMind, e.g., variable reasoning and code optimization reasoning. Furthermore, we want to augment CodeMind with a benchmark that can challenge LLMs’ code reasoning to a greater extent compared to the existing benchmarks.

### 21 Feb 2024 v3

@@ Abstract
Our extensive evaluation of nine LLMs across five benchmarks in two different programming languages using CodeMind shows that LLMs fairly follow control flow constructs and, in general, explain how inputs evolve to output, specifically for simple programs and the ones they can correctly synthesize. However, their performance drops for code with higher complexity, non-trivial logical and arithmetic operators, non-primitive types, and API calls. Furthermore, we observe that, while correlated, specification reasoning (essential for code synthesis) does not imply execution reasoning (essential for broader programming tasks such as testing and debugging): ranking LLMs based on test passing can be different compared to code reasoning [^1] The reasoning of LLMs and humans exhibit fundamental differences due to distinct nature of their cognitive processes. Our conclusions on the extent of code reasoning abilities of LLMs do not imply human-like reasoning.
cl144@illinois.edu      shizhuo2@illinois.edu      reyhaneh@illinois.edu
@@ 1 Introduction
Using CodeMind, we performed a large-scale ground theory study to assess LLMs for code reasoning. We selected nine models, including both general-purpose and Code LLMs and prompted them for IER, DER, and SR tasks on 5395 programs written in Java and Python .These programs are from five programming benchmarks, namely HumanEval (Chen et al., 2021), MBPP (Odena et al., 2021), CRUXEval (Gu et al., 2024) CodeNet (Puri et al., 2021), and Avatar (Ahmad et al., 2021). We observe that:
@@
Our contributions are (1) CodeMind framework for code reasoning that formally defines three inductive code reasoning tasks. CodeMind is open-source (CodeMind, 2024) and accepts contributions from researchers to integrate more code reasoning tasks into it; (2) a large-scale ground-theory evaluation of LLMs for code reasoning using CodeMind; and (3) a comprehensive, in-depth analysis of results that offers a catalog of root causes negatively impacting the abilities of LLMs for code reasoning. This catalog would be a valuable guideline for developing better benchmarks that truly evaluate the programming abilities of LLMs.
@@ 2 Related Work
Zhang et al. (2024) investigate transformers’ ability to learn or infer the recursive patterns from input and output pairs. They conclude that due to the inherent limitations of transformers, they may fail to learn recursion and instead find shortcut algorithms to reason about how outputs are related to inputs. Compared to this work, we evaluate LLMs regardless of architecture and training data but from the program perspective. We show LLMs can follow recursion but usually lose track of data flow due to the inability to correctly reason about loop conditions.
@@ 3 CodeMind
Program specification defines a function $S:S_{I}\rightarrow S_{O}$, where $S_{I}$ is a set of all possible inputs to the program and $S_{O}$ is a set of corresponding outputs. A code synthesized based on the implementation is usually a function $C:C_{I}\rightarrow C_{O}$. We define a program to be correct with respect to specification, if it satisfies all the following conditions:
@@ 3.1 Execution Reasoning
IER evaluates LLMs for any arbitrary code for general inductive code reasoning, which requires knowing code constructs, arithmetic and logic operations, and control flow. However, even for human developers, reasoning about their developed code is easier than any arbitrary code. Furthermore, as a self-consistency (Min et al., 2023) measurement, LLMs should be able to reason about the code they can correctly synthesize. This demands to have the following execution reasoning task.
@@ 5 LLM Evaluation on IER
  • Moving down in the table, the models face greater challenges in IER, i.e., reasoning about the execution on CodeNet and Avatar programs, compared to MBPP, HumanEval, and CRUXEval. One potential reason is the complexity of such programs as demonstrated in Figure 2. A detailed breakdown of the model’s performance (Figure 3) shows a strong negative correlation Spearman’s Rank Order Correlation (ROC) (Spearman, 1961) between CC and CRR. between CRR and CC, confirming that models struggle more in IER for a more complex code. At the same time, some models, namely Llama 2, CodeLlama, MagicCoder, StarCoder, and WizardCoder, achieve a lower performance on CRUXEval compared to HumanEval, which are less complex regarding both LoC and CC. This entails a further better understanding of what factors other than CC impact the CRR performance of the models (section 8).
@@ 6 LLM Evaluation on DER
We seek to address the critical question of how effectively the model can correctly reason about the correct programs it has generated. This evaluation requires us to align code synthesis and code reasoning tasks together. Our pipeline for evaluating DER consists of three steps: (1) following the best practices, we prompted subject LLMs for code synthesis; (2) we ran the synthesized program against existing tests; and (3) for the programs with test pass, we prompted the model for code execution reasoning using the chosen test input and under CoT style. Note that we also removed the comments from the synthesized code for fairness. We excluded the programs from CRUXEval, CodeNet, and Avatar, since these datasets are not designed for code synthesis and lack proper program specifications. Also, we could not reproduce the code synthesis results of Llama 2 and excluded that from subject LLMs. Similar to IER experiments, we set the temperature to zero to account for the non-determinism and reproducibility of the results. As a result of this design decision, our synthesis results might be different from existing leaderboards.
@@ 8 In-Depth Analysis of Results
We further analyzed the IER results, which evaluate the general ability of LLMs in code reasoning. In the first step, we wanted to see if LLMs know how different code constructs work. Without knowing the logic of each code construct, reasoning about code execution is impossible. To that end, we tagged each of $5395$ programs based on code constructs used in their implementation with the following labels: For, While, If, Try, Switch, Nested Loop, Nested If, and Basic. A program tagged with a Basic label has no special code construct. Next, we clustered the programs per tag and computed the CRR of LLMs for each cluster. Figure 5 shows the results of this analysis for the top five best-performing LLMs. We can observe that models handle conditional statements better than recursion, except for Try-Catch or Try-Except statements. Furthermore, when it comes to nested constructs, the CRR values notably drop. Impact of Loop Properties. Given that models struggle the most with recurring constructs, we focused the programs with For, While, and Nested Loop tags at the next step. We hypothesize this struggle is due to the loop’s length or determining the loop’s length. The former questions whether it is harder for the model to track the program’s data flow as loops get longer. The latter questions the abilities of models to reason about how many times a code block should be repeated, regardless of how long the length would be.
Figure 6 plots the distribution of correct versus incorrect cases and CRR values per each loop length in Java programs. Sub-figure labels show the ROC coefficients between the loop length and CRR. We can observe a moderate to strong negative correlation between the loop length and CRR of the models, i.e., CRR decreases as the loop length increases. By manually investigating the incorrect IER cases, we also noticed that LLMs mostly failed to reason about loop conditions correctly. Without knowing loop conditions and the number of iterations, it is impossible to reason about the execution correctly. Furthermore, we found cases where, although the model could reason about loop conditions and the number of iterations, it lost track of data flow in the loop, thereby predicting output incorrectly.
@@
All models failed in IER of this relatively simple program. Looking at the CoT analysis, GPT-3.5 knows the loop condition and can reason about all the variable values in the first and second iterations. However, it stays in the loop for one more. StarCoder cannot correctly reason about loop condition and iterates the code block six times. MagicCoder CoT only explains the code (correct explanation) without communicating how it reasons about variables.
@@ 9. In third iteration, amount becomes 15 and time
Impact of if-else Properties. Our manual investigations reveal that LLMs struggle with conditional constructs for two reasons: (1) confusion about branches and (2) wrong reasoning about conditional predicates. Concerning the former, we observed that LLMs sometimes struggle to recognize conditional statements. For example, regardless of which branch to take, they always go into both “if” and “else” branches to reason about execution. However, such cases rarely happen, and in most cases, LLMs struggle to analyze the conditional predicates correctly, take the wrong branch, and predict the output incorrectly.
Impact of Math and Logic Operators. Knowing how the math and logic operators work is essential to reasoning how inputs evolve into output through execution. The observations from previous analyses also underscore their importance to reason about the program control flow. Our manual investigation of failed IER cases shows the prevalence of arithmetic and logic operators in programs, with the former dominating the number of the latter. LLMs struggled more to reason about bitwise operators (e.g., $>$, $\wedge$, >>, <<, |, and &\&&) compared to arithmetic (e.g., + and -) and comparison operators (e.g., >, <).
@@
StarCoder, while failing to predict the output, correctly knows how most operators work. For example, it can reason that the bitwise AND of n and n+1 is used to determine whether integer n is a power of 2. It can also reason that the loop condition checks if the least significant bit of n is set to 0. However, it seemingly cannot understand the meaning of the return statement (setting the least significant bit to 1), hence failing to predict the correct output.
@@
Tuples and Lists consist of multiple items, and every single one of them may change during the program execution. As a result, it is unsurprising that models struggle to track the flow of inputs through potentially different execution paths and reason about a complex output as a whole. Additionally, given that manipulation of such types involves API calls, e.g., min(), next(), charAt(), reasoning about changes requires LLMs know how the APIs work, which is an additional effort.
@@ 9 Concluding Remarks
In this paper, we discussed the necessity of code reasoning tasks as an alternative to evaluate LLMs for programming tasks. We introduced CodeMind, a framework that supports several code reasoning tasks, and used CodeMind in a large-scale grounded theory study to evaluate state-of-the-art LLMs for code reasoning. Our results demonstrate that LLMs, in general, know how code constructs work and are capable of reasoning about program specification and following how inputs evolve to output through execution. However, their ability is limited as the code becomes more complex, i.e., has more complex control- or data flow, contains non-primitive types, and invokes API calls. We also observe that specification reasoning, which is essential to generate a code from a given program specification, does not mean models can also reason about code execution. We are considering two future directions based on this work. First, we plan to add more code reasoning tasks to CodeMind, e.g., variable reasoning and code optimization reasoning. Furthermore, we want to augment CodeMind with a benchmark that can challenge LLMs’ code reasoning to a greater extent compared to the existing benchmarks.